x=-10+y
-16x-2y=-2

Answer by drk(1908)

Here are the two original equations:
(i) x = -10 + 1y
(ii) -16x - 2y = -2
---
Use substitution to solve them. Put (i) into (ii) to get
-16(-10 + y) - 2y = -2
160 -16y - 2y = -2
-18y = -162
y = 9
X = -1.